Imagine you have a job registering people to vote, and you need to register one more person before the end of the day. Based on your work thus far, you know that each person you approach has a *4% probability* of registering to vote with you. What is the *expected number of people* you need to approach before you're likely to get someone to register? Please also include the *standard deviation* around your estimate, and describe how you arrive at your answer, making sure to reference the *type of distribution* this event follows.
Answer #1

Ans:

Let X=Number of trials to get first success

X will follow geometric distribution with p=0.04

Expected number of people=1/p=1/0.04=25

Variance=(1-p)/p^2

Variance=(1-0.04)/0.04^2=600

standard deviation=sqrt(600)=24.5